# Scalable Policy Evaluation with Video World Models

### v2

@@ Abstract
Training generalist policies for robotic manipulation has shown great promise, as they enable language-conditioned, multi-task behaviors across diverse scenarios. However, evaluating these policies remains difficult because real-world testing is expensive, time-consuming, and labor-intensive. It also requires frequent environment resets and carries safety risks when deploying unproven policies on physical robots. Manually creating and populating simulation environments with assets for robotic manipulation has not addressed these issues, primarily due to the significant engineering effort required and the often substantial sim-to-real gap, both in terms of physics and rendering. In this paper, we explore the use of action-conditional video generation models as a scalable way to learn world models for policy evaluation. We demonstrate how to incorporate action conditioning into existing pre-trained video generation models. This allows leveraging internet-scale in-the-wild online videos during the pre-training stage, and alleviates the need for a large dataset of paired video-action data, which is expensive to collect for robotic manipulation. Our paper examines the effect of dataset diversity, pre-trained weight and common failure cases for the proposed evaluation pipeline.Our experiments demonstrate that, across various metrics, including policy ranking and the correlation between actual policy values and predicted policy values, these models offer a promising approach for evaluating policies without requiring real-world interactions.
@@ I Introduction
Robot learning has seen remarkable strides in the development of generalist robot manipulation policies that are capable of understanding natural language instructions and performing diverse tasks across different environments and robot embodiments [black2024pi_0, brohan2022rt, octo_2023, kim24openvla]. However, as these policies and the tasks they support grow in diversity and complexity, a fundamental challenge arises: how do we systematically evaluate the capabilities of generalist robot manipulation policies in a scalable and efficient way?
@@
Figure 1: Overview. We propose using action-conditional video prediction models as world model for scalable policy evaluation. The process involves: (a) deploying a policy into the video model to generate rollout videos conditioned on action sequences and initial frame, (b) using a VLM to judge task success or failure, and (c) computing the correlation between predicted and actual policy performance to evaluate the reliability of the approach.
@@
To this end, we propose a scalable, fully automated framework for policy evaluation that leverages video models. Our framework enables the policy to interact with an action-conditioned video prediction model [arxiv25_cosmos, nvidia_cosmos_predict2, wan2025, FangqiIRASim2024] to generate rollout videos. We then use an off-the-shelf Vision-Language-Model (VLM) [gemini25_arxiv_2025, chatgpt2025] to annotate these rollouts and assess task completion, eliminating the need for human intervention and assessment. We present the correlation between the real policy performance and estimated performance from the action-conditional video model, using both simulation and real-world environments and tasks. Also, we systematically investigate the design space of diffusion-based action-conditional video prediction models and their effectiveness in evaluating robot policies. First, we examine how expanding the demonstration dataset (from 0.3 million to 1.1 million transitions) affects performance. While larger datasets improve generalization in terms of action-following, models still fail to enforce basic physical constraints, such as object permanence and rigid-body interactions. Second, we study how pre-training on large-scale text-to-video and image-to-video datasets affects generalization. Our results demonstrate that pre-trained models yield more physically plausible predictions, underscoring the promise of large-scale pre-training for building data-driven simulators. Third, we present experimental results using different pre-trained video model backbones. Finally, we highlight common failure cases in video models that may affect the reliability of policy evaluation. Our contributions can be summarized as follows:
@@
  • We conduct an analysis of the effects dataset diversity, and pre-training of action-conditional video prediction models for policy evaluation tasks.
  • We plan to release the proposed policy evaluation pipeline, including the initial frames for both real-world and synthetic tasks, the action-conditioned model, and the VLM for task completion to facilitate further benchmarking of action-conditioned video models in policy evaluation.
Supplementary materials, including additional qualitative results and an appendix are available on the project website.
@@ II Related Work
Robot Policy Evaluation. Reliable robot policy evaluation is essential for benchmarking algorithmic progress and ensuring transparency in real-world performance. In autonomous driving, policies are commonly validated in simulation across diverse scenarios before real-world deployment [dosovitskiy2017carla, yang2023unisim]. However, simulating dynamic interactions with objects, realistic materials, and complex lighting conditions in robotic manipulation remains an open challenge. Efforts such as YCB [calli2015ycb] have standardized objects for evaluation, yet standardizing the remaining factors within an environment remains a challenge. Real-world robot challenges, including the Amazon Picking Challenge [correll2016analysis] and the DARPA Robotics Challenge [krotkov2018darpa], address this by using fixed physical test sites. However, these setups demand extensive preparation and maintenance. These challenges motivate an increasing interest in using simulation to estimate policies’ success rates. SIMPLER [li24simpler] constructs physical simulators to approximate real-world environments. They show that one can deploy the policies trained with real-world data in the simulator to estimate their performance in the real world. However, this approach has two major limitations: (1) the reliance on extensive domain randomization to close the sim-to-real gap, and (2) the need for manually curated 3D data and precise environment design. In contrast, our approach eliminates manual engineering by leveraging action-conditional video prediction models to generate simulation environments directly from policies’ observations, offering a scalable and data-driven alternative to classical simulator-based evaluation methods. AutoEval [arxiv25_autoeval] presents a system that autonomously evaluates generalist robot policies in the real world. However, real-world policy evaluation still faces several challenges, such as motor overheating and long wall-clock times. Moreover, real-world evaluations are inherently imperfect and subject to noise and variability. WorldEval [worldeval] proposed using a video model for policy evaluation; however, it requires the video model to utilize features from the action decoder as conditioning signals, which limits its applicability for evaluating new policies.
@@ III-A Evaluating Robot Policies
Figure 2: Network Architecture for Action-Conditional Video Prediction with a DiT-Based Video Model such as Cosmos-Predict2-2B [nvidia_cosmos_predict2]. We process the raw action information with Fourier feature mappings [fourier] and encode it into a latent action token. The latent action token is added to a time-embedding that enables us to make the video diffusion model action-conditional.
@@ IV-A Synthetic Setting
For controllable study , we first test our policy evaluation paradigm in a synthetic setting , where the environment dynamics are treated as ground truth . In this case, dynamics in synthetic environment are treated as ground-truth dynamics. Specifically, we ask whether the video model can predict the performance ranking of policies trained entirely on synthetic data and evaluated in the same synthetic environment.
@@ IV-B2 Effectiveness of Pretrained Weights
As shown in Table II, removing pretrained weights causes a substantial drop in video prediction quality for both in-distribution and out-of-distribution data. Models trained from scratch exhibit more frequent distortions and hallucinations compared to pretrained models.
We hypothesize that pretraining provides useful priors about physical dynamics that are difficult to learn from demonstration data alone. Without these priors, models trained from scratch often display hallucination, deformation, and occlusion when the gripper interacts with objects or when objects collide (see Figure 7). This suggests that the model without pre-training lacks a robust understanding of physics, limiting its effectiveness for policy evaluation.
Policy evaluation results, summarized in Figure 3 and Table I, confirm that pretrained priors significantly improve evaluation accuracy.
@@ IV-D Failure Case Analysis
We have demonstrated the potential and feasibility of using video prediction models for policy evaluation. In this section, we discuss the limitations of our approach and highlight common failure modes observed in both simulation and real-world experiments. More quantitative analysis about failure case and capacity of the based model are provided in appendix.
@@ V Conclusion
We investigated how to adapt pre-trained video prediction models into action-conditional world models for scalable robot policy evaluation. By conditioning large video models on actions, our approach enables policy assessment without the need for costly real-world rollouts or heavily engineered simulation environments. Our analysis underscores three key factors for reliable evaluation: the value of diverse training datasets, the benefits of pretrained weights, and the failure modes that can arise during prediction.
Across both synthetic and real-world experiments, action-conditional video models accurately predict policy performance, achieving strong correlations with ground-truth success rates and supporting policy ranking. These results point to video-based world models as a promising foundation for data-driven evaluation pipelines—reducing reliance on expensive robotic infrastructure and accelerating progress toward generalist robotic policies.

### v3

@@ Abstract
Training generalist policies for robotic manipulation has shown great promise, as they enable language-conditioned, multi-task behaviors across diverse scenarios. However, evaluating these policies remains difficult because real-world testing is expensive, time-consuming, and labor-intensive. It also requires frequent environment resets and carries safety risks when deploying unproven policies on physical robots. Manually creating and populating simulation environments with assets for robotic manipulation has not addressed these issues, primarily due to the significant engineering effort required and the substantial sim-to-real gap, both in terms of physics and rendering. In this paper, we explore the use of action-conditional video generation models as a scalable way to learn world models for policy evaluation. We demonstrate how to incorporate action conditioning into existing pre-trained video generation models. This allows leveraging internet-scale in-the-wild online videos during the pre-training stage and alleviates the need for a large dataset of paired video-action data, which is expensive to collect for robotic manipulation. Our paper examines the effect of dataset diversity, pre-trained weights, and common failure cases for the proposed evaluation pipeline. Our experiments demonstrate that, across various metrics, including policy ranking and the correlation between actual policy values and predicted policy values, these models offer a promising approach for evaluating policies without requiring real-world interactions.
@@ I Introduction
Robot learning has seen remarkable strides in the development of generalist robot manipulation policies that are capable of understanding natural language instructions and performing diverse tasks across different environments and robot embodiments [black2024pi_0, brohan2022rt, octo_2023, kim24openvla]. However, as these policies and the tasks they support grow in diversity and complexity, a fundamental challenge arises: how can we systematically evaluate the capabilities of generalist robot manipulation policies in a scalable and efficient way?
@@
Figure 1: Overview. We propose using action-conditional video prediction models as world models for scalable policy evaluation. The process involves: (a) deploying a policy into the video model to generate rollout videos conditioned on action sequences and initial frame, (b) using a VLM to judge task success or failure, and (c) computing the correlation between predicted and actual policy performance to evaluate the reliability of the approach.
@@
To this end, we propose a scalable, fully automated framework for policy evaluation that leverages video models. Our framework enables the policy to interact with an action-conditioned video prediction model [arxiv25_cosmos, nvidia_cosmos_predict2, wan2025, FangqiIRASim2024] to generate rollout videos. We then use an off-the-shelf Vision-Language Model (VLM) [gemini25_arxiv_2025, chatgpt2025] to annotate these rollouts and assess task completion, eliminating the need for human intervention and assessment. We present the correlation between real policy performance and estimated performance from the action-conditional video model, using both simulation and real-world environments and tasks. Also, we systematically investigate the design space of diffusion-based action-conditional video prediction models and their effectiveness in evaluating robot policies. First, we examine how expanding the demonstration dataset (from 0.3 million to 1.1 million transitions) affects performance. While larger datasets improve generalization in terms of action-following, models still fail to enforce basic physical constraints, such as object permanence and rigid-body interactions. Second, we study how pre-training on large-scale text-to-video and image-to-video datasets affects generalization. Our results demonstrate that pre-trained models yield more physically plausible predictions, underscoring the promise of large-scale pre-training for building data-driven simulators. Third, we present experimental results using different pre-trained video model backbones. Finally, we highlight common failure cases in video models that may affect the reliability of policy evaluation. Our contributions can be summarized as follows:
@@
  • We conduct an analysis of the effects of dataset diversity and pre-training of action-conditional video prediction models for policy evaluation tasks.
  • We plan to release the proposed policy evaluation pipeline, including the initial frames for both real-world and synthetic tasks, the action-conditioned model, and the VLM for task completion, to facilitate further benchmarking of action-conditioned video models in policy evaluation.
Supplementary materials, including additional qualitative results and an appendix, are available on the project website.
@@ II Related Work
Robot Policy Evaluation. Reliable robot policy evaluation is essential for benchmarking algorithmic progress and ensuring transparency in real-world performance. In autonomous driving, policies are commonly validated in simulation across diverse scenarios before real-world deployment [dosovitskiy2017carla, yang2023unisim]. However, simulating dynamic interactions with objects, realistic materials, and complex lighting conditions in robotic manipulation remains an open challenge. Efforts such as YCB [calli2015ycb] have standardized objects for evaluation, yet standardizing the remaining factors within an environment remains a challenge. Real-world robot challenges, including the Amazon Picking Challenge [correll2016analysis] and the DARPA Robotics Challenge [krotkov2018darpa], address this by using fixed physical test sites. However, these setups demand extensive preparation and maintenance. These challenges motivate increasing interest in using simulation to estimate policies’ success rates. SIMPLER [li24simpler] constructs physical simulators to approximate real-world environments. They show that one can deploy the policies trained with real-world data in the simulator to estimate their performance in the real world. However, this approach has two major limitations: (1) the reliance on extensive domain randomization to close the sim-to-real gap, and (2) the need for manually curated 3D data and precise environment design. In contrast, our approach eliminates manual engineering by leveraging action-conditional video prediction models to generate simulation environments directly from policies’ observations, offering a scalable and data-driven alternative to classical simulator-based evaluation methods. AutoEval [arxiv25_autoeval] presents a system that autonomously evaluates generalist robot policies in the real world. However, real-world policy evaluation still faces several challenges, such as motor overheating and long wall-clock times. Moreover, real-world evaluations are inherently imperfect and subject to noise and variability. WorldEval [worldeval] proposed using a video model for policy evaluation; however, it requires the video model to utilize features from the action decoder as conditioning signals, which limits its applicability for evaluating new policies.
@@ III-A Evaluating Robot Policies
Figure 2: Network Architecture for Action-Conditional Video Prediction with a DiT-Based Video Model such as Cosmos-Predict2-2B [nvidia_cosmos_predict2]. We process the raw action information with Fourier feature mappings [fourier] and encode it into a latent action token. The latent action token is added to a time-embedding, which enables us to make the video diffusion model action-conditional.
@@ IV-A Synthetic Setting
For a controllable study, we first test our policy evaluation paradigm in a synthetic setting, where the environment dynamics are treated as ground truth. In this case, the dynamics in the synthetic environment are treated as ground-truth dynamics. Specifically, we ask whether the video model can predict the performance ranking of policies trained entirely on synthetic data and evaluated in the same synthetic environment.
@@ IV-B2 Effectiveness of Pretrained Weights
As shown in Table II, removing pre-trained weights causes a substantial drop in video prediction quality for both in-distribution and out-of-distribution data. Models trained from scratch exhibit more frequent distortions and hallucinations compared to pre-trained models.
We hypothesize that pretraining provides useful priors about physical dynamics that are difficult to learn from demonstration data alone. Without these priors, models trained from scratch often display hallucination, deformation, and occlusion when the gripper interacts with objects or when objects collide (see Figure 7). This suggests that the model without pretraining lacks a robust understanding of physics, limiting its effectiveness for policy evaluation.
Policy evaluation results, summarized in Figure 3 and Table I, confirm that pre-trained priors significantly improve evaluation accuracy.
@@ IV-D Failure Case Analysis
We have demonstrated the potential and feasibility of using video prediction models for policy evaluation. In this section, we discuss the limitations of our approach and highlight common failure modes observed in both simulation and real-world experiments. More quantitative analysis about failure cases and the capacity of the base model is provided in the appendix.
@@ V Conclusion
We investigated how to adapt pre-trained video prediction models into action-conditional world models for scalable robot policy evaluation. By conditioning large video models on actions, our approach enables policy assessment without the need for costly real-world rollouts or heavily engineered simulation environments. Our analysis underscores three key factors for reliable evaluation: the value of diverse training datasets, the benefits of pre-trained weights, and the failure modes that can arise during prediction.
Across both synthetic and real-world experiments, action-conditional video models accurately predict policy performance, achieving strong correlations with ground-truth success rates and supporting policy ranking. These results point to video-based world models as a promising foundation for data-driven evaluation pipelines, reducing reliance on expensive robotic infrastructure and accelerating progress toward generalist robotic policies.